OpenStudy (anonymous):

find percent notation 9/10

OpenStudy (anonymous):

9 * 10 /10 * 10 = ....?

OpenStudy (anonymous):

so to find the percent notation i just need to multiply 9*10

OpenStudy (anonymous):

so the answer would be 90%

OpenStudy (anonymous):

looks right to me :) To get percent notation, you need to convert whatever you have into something that is like x/100. Then that "x" is the percentage... the part out of 100%. Like on a test with 50 questions, if you miss 1 out of 50, your score is 49/50. To get a percentage, you need to multiply 49/50 by something that will make the bottom 100... but you have to multiply top & bottom by the same thing. 49/50 x 2/2 = 98/100 = 98% 24 out of 25 on a test... 24/25... 24/25 x 4/4 = 96/100 = 96%

OpenStudy (anonymous):

so it would be 9/10 = 9/100 x 10/100 = 9x10=90 10x10=100

OpenStudy (anonymous):

close... 9/10 = 9/10 x 10/10 = (9 x 10) / (10 x 10) = 90/100 = 90%

OpenStudy (anonymous):

the term you multiply by your fraction has to be equal "1"... you can't just multiply by any number... so you choose a fraction like (10/10) or (2/2) or whatever you need to transform your ORIGINAL denominator into 100. In this case, you had 9/10... so you choose (10/10) If you had a fraction like 48/50, you need to use (2/2) to get 48x2 / 50x2 = 96/100

OpenStudy (anonymous):

ok i understand thank you so very much

OpenStudy (anonymous):

so for 7/8 how do i know what numbers to use to get the %

OpenStudy (anonymous):

yes, this one is harder :) well, you could do it like a ratio... 7/8 = x/100 And then cross multiply to get x which is your percentage.

OpenStudy (anonymous):

What did you get for 7/8 as a percentage?

OpenStudy (anonymous):

87

OpenStudy (anonymous):

you rounded it, it looks like... 7/8 = x/100 cross multiply: 700 = 8x Solve for x = 700/8 = 0.875 = 87.5%

OpenStudy (anonymous):

i did 7/8 =x/100 then i crossed multiply to get 800 then divided by 7 to get 87 i just didnt include the remainder

OpenStudy (anonymous):

do i need to include the remainder

OpenStudy (anonymous):

yes, you wouldn't want a remainder, but if you did that part on a calculator, you would get 87.5%

OpenStudy (anonymous):

hope this helped... I have to leave, but if you have other questions, post them as new questions and someone else will help you. Good luck!

OpenStudy (anonymous):

thank you